OpenStudy (anonymous):

Find an orthonormal basis for the subspace S = {(x1, x2, x3) : x1 + 2x2 − x3 = 0} ⊂ R^3

OpenStudy (zarkon):

[-1,1,1] [1,0,1]

OpenStudy (phi):

hmm, did I miss something?

OpenStudy (zarkon):

that is just orthogonal..you need to normalize it

OpenStudy (zarkon):

\[x_1+2x_2-x_3=0\] is just a plane in \[R^3\] find two vectors in the plane that are orthonormal.

OpenStudy (phi):

thanks

OpenStudy (phi):

Definitely got that backwards. What we have is \[[1 -2 -1]\left[\begin{matrix}x1 \\ x2 \\x3\end{matrix}\right]=0\] Now find the basis for the null space of [1 -2 1], and normalize, which is what Zarkon did.

OpenStudy (anonymous):

Yeah I understand it better when you rewrote it phi, how do you normalize it then...

OpenStudy (zarkon):

\[\|<a,b,c>\|=\sqrt{a^2+b^2+c^2}\] so our unit vector in the sam direction of \[<a,b,c>\] is \[\frac{<a,b,c>}{\|<a,b,c>\|}\]

OpenStudy (zarkon):

'sam' should be same